OpenStudy (anonymous):

write the expression as a single trigonometric function..... sin 9x cos 3x - cos 9x sin 3x

OpenStudy (anonymous):

Here, \(\sin9x \cos 3x - \cos 9x \sin 3x\) Let 9x = a and 3x = b \(\sin a \cos b - \cos a \sin b\)

OpenStudy (anonymous):

That is an identity : \(\sin (a-b) = \sin a \cos b - \cos a \sin b\)

OpenStudy (anonymous):

So we have : \(\sin 9x \cos 3x - \cos 9x \sin 3x = \sin (9x-3x)\) = ? Can you tell me now @kausarsalley ?

Parth (parthkohli):

@kausarsalley ...?

OpenStudy (anonymous):

\[\sin (9x -3x)= \sin 9x \cos 3x - \cos 9x \sin 3x\]

Parth (parthkohli):

Yes, but can you simplify \(\sin(9x - 3x)\)?

OpenStudy (anonymous):

no please.....

OpenStudy (anonymous):

Can you simplify 9x-3x?

OpenStudy (anonymous):

\(\color{blue}{9x-3x=?}\)

OpenStudy (anonymous):

\[9x - 3x = 6x\]

OpenStudy (anonymous):

Yes so \(\sin (9x-3x) = ?\)

OpenStudy (anonymous):

sin 6x..

Parth (parthkohli):

\[\checkmark\]

OpenStudy (anonymous):

That's it. great work.

OpenStudy (anonymous):

thank you both alot! :)

Parth (parthkohli):

Not me, not me.

OpenStudy (anonymous):

You're welcome.